• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 17:27:07 on October13,2007

    Entry number 208919

    Followups:

    keyword=Recabling of S2

    Scalers:



    1) I took the two S2m ribbon cables for the left and right PMTs plugged

        into the 880 ns and 870 ns PEC delay boxes. The right outputs from these

        delay boxes were used for the S2m scalers and plugged into the module

        named 'cea0'.



    2) I took the "Left S2" scaler cable and plugged it into the 4518 delay

        module for S2, which plugs into channels 1-16 of 'cea0'.

        Added an additional label: "S2-L/R Scaler Oct 08, 2008".



    3) I took the "Right S2" scaler cable and plugged it into the 4516 logic

        module for S2, which plugs into channels 17-32 of 'cea0'.

        Added an additional label: "S2-L/R Scaler LandR Oct 08, 2008".



    Hence, channels 1-6 contain the S2-L signals and channels 9-14 contain the

    S2-R signals. Then channels 17-22 contain the (Left PMT .and. Right PMT)

    signals from the paddles.



    In the scaler.map file, the scalers for S2m were from crate #7, slot #1,

    start chan 0, and nchan 32. These same channels are now being used for S2

    as defined above but of course, offset by -1.



    TDCs:



    1) For the S2 TDCs, I am using the 870 ns PEC delay #2 box that was used

        for the S2m right PMTs. I took an unused ribbon cable labeled "VME CEB1

        1" to connect the 4518 second output to the PEC delay #2 box input. The output

         cable goes downstairs to ROC2 in slot #16 (TDC), channels 16-31. Added an

        additional label: "S2-L/R TDC Out Oct 08, 2008".



    2) Hence, channels 16-21 are for the S2-L signals and channels 24-29 are

        for the S2-R signals.



    3) As a reference from the S2m database file, the TDCs signals were connected

        to crate #2, slot #16, channels 0-15 for the left PMTs and channels 16-31 for

        the right PMTs.



    ADCs:



    1) The signals from the 4413 discrim. plug into the patch panel below the

        trigger crate into row B. This row was previously used for S2m signals

        S2-1L through S2-12L. The BNC connector 'B4' was not used for S2-4L, it

        was instead plugged into 'A13'. I assume this means there is a problem

        with 'B4', since nothing is plugged into it. I borrowed an unused analogue

        ribbon cable from the aerogel detector. Added an additional label to analogue

        ribbon cable: "S2-L/R ADC Oct 08, 2008".



    2) The signals from this patch panel go downstairs to a similar patch

        panel. I verified that row B was being used for S2m signals. Again

        connector 'B4' is not in use, and the signal for S2-4L was connected to

        'A13'.



    3) The signals then plug into ROC1, slot #24 (ADC), channels 48-63 via an

        analogue ribbon cable. Added an additional label to analogue ribbon cable:

        "S2-L/R ADC Oct 08, 2008".



    4) Hence, channels 48-53 are for S2-L signals and 54-59 are for S2-R signals.